OpenStudy (anonymous):

Larry's Landscaping is building a circular flower bed that can have a maximum area of 100 square feet. To the nearest foot, the longest radius that Larry's Landscaping can use to make the circle is (blank) feet.

OpenStudy (anonymous):

Lets see if we can write an inequality equation to represent the statement: The area of the circle has to be less than or equal to 100

OpenStudy (anonymous):

Circle area <= 100

OpenStudy (anonymous):

area of circle (3.14r^2) < 100.

OpenStudy (anonymous):

Exactly. Use the less than or equal to symbol. Now , can you solve that for r?

OpenStudy (anonymous):

What do I subtract in order to figure out this equation??

OpenStudy (anonymous):

you don't subtract. first, we want to get rid of that 3.14. to do that, divide both sides by 3.14

OpenStudy (anonymous):

(1/3.14) 3.14r^2 <=100 (1/3.14) r^2 <= 100 Like this??

OpenStudy (anonymous):

I meant 100/3.14, which would be 31.84

OpenStudy (anonymous):

yes... so you have r^2 <=31.86 now we want to turn that r^2 into an r. to do that, take the square root of both sides.

OpenStudy (anonymous):

I ended up with 5.64. So, to the nearest foot that would be about 6 feet. Would this be correct?

OpenStudy (anonymous):

that's it! that's what i got.

OpenStudy (anonymous):

Alright, thank you very much!!

OpenStudy (anonymous):

no problem. good luck